If the numerator of a fraction is increased by 200% and the denominator is increased by 300% the
Resultant fraction is 6/11. What was the original fraction?

Answer by dkppathak(439)

let fraction is x/y
when numerator increase by 200%
means numerator will be x+2x =3x
and denominator increased by 300% means denominator will be y = 3y =4y
new fraction will be 3x/4y =6/66 (given)
we can say
3x/4y = 6/11
dividing numerator by 3 and denominator by 4 both side we will get
x/y =6/3/11/4
x/y=6x4/11x3
x/y =24/33
x/y =8/11
answer
fraction will be 8/11